OpenStudy (anonymous):

Find the slope of the line that passes through these points. (5/2,-1),(1/2,9)

OpenStudy (anonymous):

the slope of a line passing through (x1,y1) and (x2,y2) is m=(y2-y1)/(x2-x1)

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

slope, m =(9-(-1))/(1/2-5/2) m=10/(-2) m=-5